# Evaluating Contextualized Representations of (Spanish) Ambiguous Words: A New Lexical Resource and Empirical Analysis

### v2

@@ 1 Introduction
We first provide a survey of related work on ambiguity and interpretability in LMs (Section 2), then present a novel dataset of human relatedness judgments about Spanish ambiguous words—in context (SAW-C) (Section 3). Section 4 documents our use of this dataset to empirically probe the representation of ambiguous words in pre-trained spanish LMs, focusing first on pre-registered analyses [^2] A link to the pre-registration, code, and data will be provided after the anonymity period has passed. of the monolingual BETO (Cañete et al., 2020). We next systematically compare multiple monolingual Spanish and multilingual BERT-based models (Section 5). We conclude with a discussion of the implications of the current results (Section 6), as well as limitations and directions for future work (Section 7).
@@ 2 Related work
Using more controlled sentence stimulus design, coupled with empirically collected human benchmarks, prior studies in English have shown that BERT-based models’ contextualized embeddings capture some—though not all—variance in human similarity (Haber and Poesio, 2020) and relatedness (Trott and Bergen, 2021; DeLong et al., 2023) judgments for ambiguous English words. Some of this work has also argued that the continuous nature of LM contextualized representations makes them well-suited as models of human word meanings, which are likely graded to some extent Elman (2009); Li and Armstrong (2024); Li and Joanisse (2021); Trott and Bergen (2023); Rodd et al. (2004); Nair et al. (2020)—though importantly, may also exhibit marked categoriality compared to LM representational spaces Trott and Bergen (2023).
@@ 3.1 Materials
All sentences were developed by two native Spanish speakers. $102$ target words were drawn from noun lists collected in earlier studies of Spanish lexical ambiguity (Estévez Monzó, 1991; Fraga et al., 2017), as well as spontaneously generated and then verified using the online Real Academia Española [^3] https://www.rae.es/ dictionary. We excluded wordform meanings that corresponded to distinct parts of speech, but we accepted a small fraction ($8/102$) of nouns whose grammatical gender changed across meanings. Each target ambiguous noun was embedded within sentence pairs that differed by a single modifier [^4] For $187/812$ sentence pairs, the modifiers varied along more than just a single word, as was the case when the modifiers required different prepositions, or the grammatical gender of the modifier differed across sentences and required distinct determiners and contractions. Of these, $173$ pairs differed by $1$ word; $14$ differed by $2$ words., termed context cue [^5] Context cues were adjectival modifiers for $100/102$ target nouns; verbs for $2/102$.. The average number of words in sentences was $4.72$. The context cues across the sentence pair could evoke either the Same or Different sense of the word across the contexts.
@@ 3.3 Procedure
Each sentence pair was presented on a separate page. Participants were instructed to rate each word on a scale from 1 (totalmente sin relación, “totally unrelated”) to 5 (mismo significado, “same meaning”). The target word (e.g., aceite, “oil”) was centered in larger font, and the target sentences were presented side-by-side (the side was randomized across trials). Participants indicated their response via button-press; see Figure 1 for an example. We also included one “catch” trial in the experiment, which simply contained the same sentence, repeated (i.e., the correct answer was $5$).
@@ 6 Discussion
Lastly, this work contributes to expanding the linguistic diversity of both human-annotated benchmarks and interpretability research. Although Spanish ranks among the most widely spoken languages, it suffers from a surprising dearth of resources, pre-trained models, and interpretability research—particularly when compared to English. In one study, a sample of 550 corpora (spanning 22 languages) contained >$50\%$ English-language corpora, while <$10\%$ represented Spanish-language corpora Anand et al. (2020). A wider research perspective—considering varied languages—is critical for ensuring the generalizability of findings.
@@ 7.1 Limitations of the Dataset
SAW-C is limited in scope, containing only $812$ sentence pairs. This is considerably smaller than English benchmarks such as BLiMP Warstadt et al. (2020), which contain tens of thousands of examples. However, it is larger than or comparable in size to other, more targeted datasets involving crowd-sourced human annotations Erk et al. (2013); Haber and Poesio (2020); Trott and Bergen (2021). Relative to specifically Spanish-language datasets, ours is the only one we are aware of that collects human judgments for target ambiguous words embedded within minimal pair stimuli [^11] For examples of other Spanish-language datasets that collect human judgments under different experimental conditions, see Estévez Monzó (1991); Gómez-Veiga et al. (2010); Domínguez et al. (2001); Haro et al. (2017); Fraga et al. (2017).. Importantly, SAW-C includes not only the sentence pairs but also over $10000$ validated human judgments from $131$ participants about those sentence pairs.
Another limitation concerns stimulus generation. The ambiguous nouns included in our dataset were spontaneously produced by native (Caribbean) Spanish speakers, or selected from previously published lists (Estévez Monzó, 1991; Fraga et al., 2017), rather than via automated searches. Spontaneous production of ambiguous words may (1) overrepresent homonymous words (Estévez Monzó, 1991), and (2) underrepresent words whose multiple meanings have large dominance asymmetries Duffy et al. (1988). In future work, we intend to collect dominance norms for these items. Differences in Spanish varieties may have also led some items to appear awkward to annotators from different nationalities. We address this concern by showing strong correlations in relatedness judgments from the three heavily represented groups (Chile, Spain, and México, Section 3.4), but other groups’ relatedness judgments may not be represented in this sample.
@@ A.1 LLM Specifications and Dataset Tokenization
Table 2: Spanish language model properties and training procedures. Models are cased (distinguish between upper and lowercase characters) unless otherwise specified. All monolingual models are trained on Spanish-language corpora; multilingual models include Spanish-language corpora. Model Notes: For mBERT, (a) the corpus size per language varied and we are unsure of the total corpus size12, (b) it was unclear to us whether the current version of the model on HuggingFace is updated with the whole-word masking (WWM) technique13 during pre-training. For RoBERTa-BNE models, it was unclear to us whether authors used dynamic masked (DM) modeling, as in the English RoBERTa. For XLM-RoBERTa, the corpus size per language may have varied, but we were uncertain to what extent pretraining text was sampled proportionally to its representation in the corpus (Conneau et al., 2020). Acronyms: BNE: Biblioteca Nacional de España (National Library of Spain); byte-BPE: byte-level Byte-Pair Encoding; DistilLoss: Distillation loss (Sahn et al., 2019; Cañete et al., 2022); DM: Dynamic Masking (Liu et al., 2020); MLM: Masked Language Modeling; NSP: Next Sentence Prediction; SP: SentencePiece; WWM: Whole-Word Masking; WP: WordPiece. This summary represents our best attempt at gathering and reconstructing some model specifications—they are necessarily incomplete, and may contain inaccuracies borne from either a lack of knowledge regarding more recent updates or an imperfect understanding of the training protocols as described in the relevant primary literature and repositories.
1212footnotetext: See https://github.com/google-research/bert/blob/master/multilingual.md for available details on language-specific corpus size selection.1313footnotetext: See https://github.com/google-research/bert for notes on the English BERT pre-training update using whole-word masking.

### v3

@@ 1 Introduction
We first provide a survey of related work on ambiguity and interpretability in LMs (Section 2), then present a novel dataset of human relatedness judgments about Spanish ambiguous words—in context (SAW-C) (Section 3). Section 4 documents our use of this dataset to empirically probe the representation of ambiguous words in pre-trained spanish LMs, focusing first on pre-registered analyses [^2] Code and data to run all analyses are available at: https://github.com/seantrott/spanish_norms. Preregistration is available at: https://osf.io/n5htp. of the monolingual BETO (Cañete et al., 2020). We next systematically compare multiple monolingual Spanish and multilingual BERT-based models (Section 5). We conclude with a discussion of the implications of the current results (Section 6), as well as limitations and directions for future work (Section 7).
@@ 2 Related work
Using more controlled sentence stimulus design, coupled with empirically collected human benchmarks, prior studies in English have shown that BERT-based models’ contextualized embeddings capture some—though not all—variance in human similarity (Haber and Poesio, 2020) and relatedness (Trott and Bergen, 2021; DeLong et al., 2023) judgments for ambiguous English words. Some of this work has also argued that the continuous nature of LM contextualized representations makes them well-suited as models of human word meanings, which are likely graded to some extent Elman (2009); Li and Armstrong (2024); Li and Joanisse (2021); Trott and Bergen (2023); Rodd et al. (2004); Nair et al. (2020); Zamora-Reina et al. (2022)—though importantly, may also exhibit marked categoriality compared to LM representational spaces Trott and Bergen (2023).
@@ 3.1 Materials
All sentences were developed by two native (Puerto Rican) Spanish speakers, both authors in this study. $102$ target words were drawn from noun lists collected in earlier studies of Spanish lexical ambiguity (Estévez Monzó, 1991; Fraga et al., 2017), as well as spontaneously generated and then verified using the online Real Academia Española [^3] https://www.rae.es/ dictionary. We excluded wordform meanings that corresponded to distinct parts of speech, but we accepted a small fraction ($8/102$) of nouns whose grammatical gender changed across meanings. Each target ambiguous noun was embedded within sentence pairs that differed by a single modifier [^4] For $187/812$ sentence pairs, the modifiers varied along more than just a single word, as was the case when the modifiers required different prepositions, or the grammatical gender of the modifier differed across sentences and required distinct determiners and contractions. Of these, $173$ pairs differed by $1$ word; $14$ differed by $2$ words., termed context cue [^5] Context cues were adjectival modifiers for $100/102$ target nouns; verbs for $2/102$.. The average number of words in sentences was $4.72$. The context cues across the sentence pair could evoke either the Same or Different sense of the word across the contexts.
@@ 3.3 Procedure
Each sentence pair was presented on a separate page. Participants were instructed to rate each word on a scale from 1 (totalmente sin relación, “totally unrelated”) to 5 (mismo significado, “same meaning”) [^7] For alternative ordinal annotation schemes in relatedness datasets, see Schlechtweg et al. (2018, 2024, 2025). The target word (e.g., aceite, “oil”) was centered in larger font, and the target sentences were presented side-by-side (the side was randomized across trials). Participants indicated their response via button-press; see Figure 1 for an example. We also included one “catch” trial in the experiment, which simply contained the same sentence, repeated (i.e., the correct answer was $5$).
@@ 6 Discussion
Lastly, this work contributes to expanding the linguistic diversity of both human-annotated benchmarks and interpretability research. Although Spanish ranks among the most widely spoken languages, it suffers from a surprising dearth of resources (with notable exceptions, Baldissin et al. (2022)), pre-trained models, and interpretability research—particularly when compared to English. In one study, a sample of 550 corpora (spanning 22 languages) contained >$50\%$ English-language corpora, while <$10\%$ represented Spanish-language corpora Anand et al. (2020). A wider research perspective—considering varied languages—is critical for ensuring the generalizability of findings.
@@ 7.1 Limitations of the Dataset
SAW-C is limited in scope, containing only $812$ sentence pairs. This is considerably smaller than English benchmarks such as BLiMP Warstadt et al. (2020), which contain tens of thousands of examples. However, it is larger than or comparable in size to other, more targeted datasets involving crowd-sourced human annotations Erk et al. (2013); Haber and Poesio (2020); Trott and Bergen (2021). Relative to specifically Spanish-language datasets, ours is the only one we are aware of that collects human judgments for target ambiguous words embedded within minimal pair stimuli [^12] For examples of other Spanish-language datasets that collect human judgments under different experimental conditions, see Estévez Monzó (1991); Gómez-Veiga et al. (2010); Domínguez et al. (2001); Haro et al. (2017); Fraga et al. (2017); Zamora-Reina et al. (2022); Baldissin et al. (2022).. Importantly, SAW-C includes not only the sentence pairs but also over $10000$ validated human judgments from $131$ participants about those sentence pairs.
Another limitation concerns stimulus generation. The ambiguous nouns included in our dataset were spontaneously produced by native (Caribbean) Spanish speakers, or selected from previously published lists (Estévez Monzó, 1991; Fraga et al., 2017), rather than via automated searches. Spontaneous production of ambiguous words may (1) overrepresent homonymous words (Estévez Monzó, 1991), and (2) underrepresent words whose multiple meanings have large dominance asymmetries Duffy et al. (1988). In future work, we intend to collect dominance norms for these items. Differences in Spanish varieties may have also led some items to appear awkward to annotators from different nationalities (Baldissin et al., 2022; Lipski, 2014). We address this concern by showing strong correlations in relatedness judgments from the three heavily represented groups (Chile, Spain, and México, Section 3.4), but other groups’ relatedness judgments may not be represented in this sample.
@@ A.1 LLM Specifications and Dataset Tokenization
Table 2: Spanish language model properties and training procedures. Models are cased (distinguish between upper and lowercase characters) unless otherwise specified. All monolingual models are trained on Spanish-language corpora; multilingual models include Spanish-language corpora. Model Notes: For mBERT, (a) the corpus size per language varied and we are unsure of the total corpus size13, (b) it was unclear to us whether the current version of the model on HuggingFace is updated with the whole-word masking (WWM) technique14 during pre-training. For RoBERTa-BNE models, it was unclear to us whether authors used dynamic masked (DM) modeling, as in the English RoBERTa. For XLM-RoBERTa, the corpus size per language may have varied, but we were uncertain to what extent pretraining text was sampled proportionally to its representation in the corpus (Conneau et al., 2020). Acronyms: BNE: Biblioteca Nacional de España (National Library of Spain); byte-BPE: byte-level Byte-Pair Encoding; DistilLoss: Distillation loss (Sahn et al., 2019; Cañete et al., 2022); DM: Dynamic Masking (Liu et al., 2020); MLM: Masked Language Modeling; NSP: Next Sentence Prediction; SP: SentencePiece; WWM: Whole-Word Masking; WP: WordPiece. This summary represents our best attempt at gathering and reconstructing some model specifications—they are necessarily incomplete, and may contain inaccuracies borne from either a lack of knowledge regarding more recent updates or an imperfect understanding of the training protocols as described in the relevant primary literature and repositories.
1313footnotetext: See https://github.com/google-research/bert/blob/master/multilingual.md for available details on language-specific corpus size selection.1414footnotetext: See https://github.com/google-research/bert for notes on the English BERT pre-training update using whole-word masking.